SkinnyMinnie20:

Use the net to find the lateral area of the prism. 432 m2 270 m2 180 m2 225 m2

SkinnyMinnie20:

1 attachment
Vocaloid:

since lateral area excludes the top and bottom bases we can simply find the area of the four middle rectangles

Vocaloid:



Vocaloid:

if area = Length*width, what is the area of this rectangle?

SkinnyMinnie20:

is it 45

Vocaloid:

good, let's keep going

Vocaloid:



Vocaloid:

area = ?

Vocaloid:

still there?

SkinnyMinnie20:

yes i am sorry

SkinnyMinnie20:

it would be 45 again

Vocaloid:

think carefully if the length is 15 and the width is 6, and area = length*width, area = 15*6 = ?

SkinnyMinnie20:

90

Vocaloid:

good now if we look at the other two rectangles, they're the same dimensions as the first two rectangles so lateral area = 90 + 45 + 45 + 90 = 270 m^2 = your answer

SkinnyMinnie20:

thank you very much